### v3

# Regular triangle unions with maximal number of sides

[…]
2025
Regular triangle unions with maximal number of sides
[…]
Intro
Latvian, Lithuanian and (now extinct) Old Prussian constitute a Baltic branch of Indoeuropean (IDE) languages. During a Baltic Culture Festival in Northern Lithuania, August 2024, the author fixed his eyes on a scene decoration, where Latvian national symbol Auseklis [^1] From Latvian aust, Lithuanian aušti, ‘to dawn’, which has common IDE roots with English East. was shown. It forms an octagonal star (Fig. 1, left), which is just a standard octagram ($\{8,3\}$-gon in Schläfli notation). Though this cross is supposed to be drawn without lifting a stick from the ground (the reason behind this restriction is rooted, of course, in magic; same picture, left, where simple and double Lietuvēns crosses are shown), one clearly can distinguish $4$ triangles whose union forms a simple $16$-gon. With few more thoughts, it is immediate how one can increase edge count by slightly deforming all triangles. It was therefore natural to wonder what is the best possible result. Such is a subjective genesis of the problem. And magic surely works! The author could not anticipate that hypostases of both crosses will eventually emerge in this purely mathematical endeavour. Yet, the earliest to rise will be $\ldots$ the Star of David. But let us go to the basics first.
Figure 1: Left – a stamp dedicated to Latvian national Front (a movement that eventually led to liberation from Soviet occupation) anniversary featuring Auseklis – a Latvian national symbol. Right – a double and simple Lietuvēns cross.
Unions
[…]
Take $n$ $\triangle$. Its edges will henceforth be called legs, while edge will be reserved solely for a $\bigcup$ polygon. Any two of $\triangle$ intersect in at most $6$ points, and it is straightforward to contrive an example with $\sim 2n^{2}$ edges on the boundary of the $\bigcup$. However, in such construction many of the $\triangle$ appear to be very slim. If all $\triangle$ happen to have angles at least $\delta>0$ (so called $\delta$-fat"), the complement of the $\bigcup$ then has at most $O_{\delta}(n)$ connected components, and the boundary consists of $O_{\delta}(n\log\log n)$ edges fat . What happens, though, if the $\bigcup$ has no holes at all and its interior is connected? Let us formulate this precisely.
Problem.
Given an integer $n\geq 1$. Suppose, a simple polygon is a $\bigcup$ of $n$ $\triangle$. What is the maximal number of sides it can have? The following also applies.
  • 1)
    No further restrictions. This defines seq. $i(n)$.
  • 2)
    Vertices of $\triangle$ along the border of the $\bigcup$ are arranged in a cyclic way – seq. $e(n)$.
  • 3)
    Vertices of $\triangle$ are on the unit circle – seq. $c(n)$.
  • 4)
    Vertices are on the unit circle and are arranged in the cyclic way – seq. $r(n)$.
It may seem surprising that the question “how many edges exactly?" has not been asked before. Yet, there are reasons for that. The seq. $i(n)$ is the one where existing techniques and results apply most. Take a $\bigcup$ of $n$ $\triangle$ with $i(n)$ edges. Find a $\triangle$ which contributes least to the common boundary. Remove it. It is not difficult to see this reduces the edge count of the outer border by at most $\frac{i(n)}{n}+3$. Thus, as if
$i(n-1)\geq i(n)-\frac{i(n)}{n}-3.$
This would imply $i(n)=O(n\log n)$. Sure, such a naïve approach is not strictly valid: we do not know whether removal of 11 $\triangle$ leaves no holes in the $\bigcup$. But a certain aspect of the phenomenon is captured. In the Section ‘Lower Envelopes’, employing the theorem of Micha Sharir Ady Wiernik (deep and unexpected result for any newcomer to the topic), we will see that one cannot expect to do much better than that: $i(n)=\Omega(n\alpha(n))$, where $\alpha(n)$ is the functional inverse of the so-called Ackermann function. It tends to $\infty$ with an argument, but extremely slowly. Its value is $3$ in the interval $[5,16]$, and $4$ in the interval
$\Bigg[17,\underbrace{2^{2^{\ldots 2^{2}}}}_{65536\text{ copies}}\,\Bigg].$
It is impossible for a sane mind to fathom a place on the line of positive integers where this function surpasses $5$. There is a handful of other reasons why a plain answer is not to be expected.
Figure 2: I=mnmn+3I=mn-m-n+3 number of intersections of $n$-gon and $m$-gon for odd $m,n$ ($m=9$, $m=7$, $I=50$ here).
However, the situation with cyclically arranged vertices is slightly different. Such $\triangle$ $\bigcup$ will be called regular. The current paper, though mostly expository and derivative, introduces such objects for the first tiome and gives a new additional tool in dealing with $\bigcup$. Namely, triangulation shifts. Theoretically, an answer becomes within a reach. Sure, there exists problems in combinatorial geometry where an arm-stretch distance is still a considerable one.
For example, it is known that two simple polygons with odd numbers ($m$ and $n$) of vertices can intersect in at least mnnm+3mn-n-m+3 points (Fig. 2), and this is conjectured to be the peak value. Ackerman et al. ackerman proved that the maximum is indeed mnmn+Cmn-m-n+C. Accordingly, this is optimal, apart from the exact value of $C$ (still not rigorously known to be equal to $3$).
Henceforth, we will concentrate on $e(n)$, referred here as the Politrigon seq.. It now has a number A375986 in the OEIS oeis . What are its initial values? For $n=1$, there is nothing to prove. The $\bigcup$ of 22 $\triangle$, if a simple $M$-gon, can have any number of sides in the range $3\leq M\leq 10$, and also $M=12$. The latter is mostly familiar from a pleasingly symmetric hexagram known as the star of David. In order to prove that $e(3)=22$, case-by-case analysis can be performed. The example ($n=1,2,3$) is presented in Fig. 11 (three brown $\triangle$).
Figure 3: Regular $\bigcup$ of $4$ triangles with $33$ sides.
Figure 4: $10$ lines forming $25$ non-overlapping $\triangle$.
Figure 5: All 44 perfect $\bigcup$ of triangles. These are exactly the configurations satisfying (n+1)e(n)=(12n6)n(n+1)e(n)=(12n-6)n. Allegorically speaking, the line of succession to the Eye of Providence continues with the Star of David, then with the Pentastar, and terminates with the Octastar. The last two have only rotational symmetry. Combinatorics of Pentastar cannot be reached from Octastar by removing 33 $\triangle$. This excludes any inductive construction of the best $\bigcup$.
Moving onto $n=4$, similar computer-assisted overhaul show that $e(4)=33$ (Fig. 3). If we knew all possible arrangements of $3n$ lines in the affine plane (seq. A090338), this catalogue would encode, among other things, the Politrigon seq. (and $i(n)$, too). Such list (for $n$ lines) would also encode the solution to the Kobon $\triangle$ problem (seq. A006066). The latter asks how many non-overlapping $\triangle$ can be formed by $n$ lines (see Fig. 4 depicting one of maximal configurations for $n=10$). In its full generality, Kobon’s problem is still unsolved. It is obvious that such full register of line configurations is of no use for large $n$ do to extreme growth (of order $2^{\Theta(n\log n)}$) of their total number. In fact, the last exact value known is for $9$ lines, which amounts to $3.111.341$ simple arrangements. In our setting, this would settle, with some effort, the case $n=3$. Clearly, in all derivatives of line arrangement problem, other method should be looked for.
And it does exist in regular $\triangle\bigcup$ case. With such tools at hand, it will be possible to demonstrate the unconditional inequality $(n+1)e(n)\leq(12n-6)n$. This leads to the discovery of two “perfect" $\triangle$ $\bigcup$ showing that $e(5)=45$ and $e(8)=80$ are possible. Together with a $\triangle$ and a hexagram, this is the complete list of maximally saturated $\triangle$ $\bigcup$. In other words, the only cases where the equality is achieved (Fig. 5). (exactly the very reincarnation of two magic crosses hinted to in the Introduction). For $n\neq 1,2,5,8$, strict inequality must hold.
Theorem..
The seq. $e(n)$ starts as $3$, $12$, $22$, $33$, $45$, $56$, $67$, $80$, $91$. It also satisfies the following:
$11\leq e(n+1)-e(n)\leq 14\quad(forn\geq 3),$
$e(n)\leq 12n-18+I(10\leq n\leq 13)\quad(forn\geq 10).$
The first inequality gives situational lower bound $e(n)\geq 11n-8$ ($n\geq 10$). This can always be improved, provided we manage to find a concrete $N\geq 10$ and a configuration of NN $\triangle$ with more than 11N811N-8 edges.
The upper bound now stands in sharp contrast with a superlinear growth of the function $i(n)$. It not just bounds $e(n)$ from above, but in fact gives the exact combinatorial bound on first approximation. Without explaining this vague phrase now, here is a little trick to be played on the readers. Among all triangulations of a regular $n$-gon, choose the one whose maximal sum of number of diagonals adjacent at three consecutive vertices is minimal. Denote this minimum by $\delta(n)$ (see Fig. 6). One of the many interpretations of Catalan numbers $C_{n}=\frac{1}{n+1}\binom{2n}{n}$ is that $C_{n-2}$ gives the number of triangulations of a convex $n$-gon. Since there are n3n-3 diagonals, $\delta(n)\geq\big\lceil 6-\frac{18}{n}\big\rceil$. The following statement is therefore a result of full check ($n\leq 17$) and explicit construction ($n\geq 18$):
Proposition 1.
The sequence $\delta(n)$, $n\geq 3$, starts as $0,2,3,3,4,5,4,5,5,5,5,5$, and then stabilizes at the value $6$ (see Fig. 6).
The reader can verify that for $n\geq 2$, $\mathscr{U}(n)=12n-12-\delta(n+1)$ coincides with exact value of the seq. $e(n)$ and gives the exact upper combinatorial bound for $n\geq 10$. The relation of $\triangle$ $\bigcup$ to triangulations of $n$-gons is unexpected, and will be clarified later. But now, let us discuss what does it mean “combinatorial" vs. “geometric". [^3] Since we started from linguistics, it is apt to emphasize here that suffix dilemma ‘-ial’ vs. ‘-ic’ essentially coincides with whether it is of Latin or of Greek origin. In other words, what are pseudolines and what is stretchability.
Figure 6: Triangulations of regular polygons which minimize the maximum of number of diagonals adjacent at three consecutive vertices. This minimum is denoted by $\alpha(n)$
Pseudolines vs. Lines
Apart from A006066 and A090338, there are other seq. in the OEIS made from kindred ingredients. This means: seq., whose definition employs only well-familiar objects from Euclidean geometry, but whose calculation requisites rather delicate arrangements of these. One of such seq. was popularized by Neil Sloane and the community of OEIS during the “ranking process" invoked to clarify which seq. should be assigned a honorary label A250.000. A runner-up in this vote, now a seq. A250.001 (first proposed by Jonathan Wild), asks to find the number of ways to draw $n$ circles in the affine plane sloane ; oeis (definitions of admissible and equivalent configurations are intuitive and can be found in the references cited). In Figure 7, three (out of $173$) arrangements of $4$ circles are reproduced.
Figure 7: Three ways to draw 4 circles. Configuration $\mathbf{c}$ has $12$ boundary arcs.
Even having a ready example in front of the eyes, it takes some practice to correctly drag around centres and radii of four circles in order to achieve configurations $\mathbf{a}$ and $\mathbf{b}$. See a recent survey cutler for many more examples of such seq., associated with cutting of plane with all kinds of “exotic knives".
On the other hand, pseudoline arrangements and stretchability problem are not only close in spirit, but play a significant role in determining $e(n)$. A pseudoline is a simple connected curve in the projective plane such that its removal does not disconnect $\mathbb{R}P^{2}$. Arrangement of pseudolines is a finite set of them so that each two meet exactly once, and there are at least two such intersection points (the affine case is modified accordingly). A survey by Felsner and Goodman felsner is another enriching reading. The most transparent arrangement of projective pseudolines that cannot be realized with lines (that is, “non-stretchable") is given by a twisted Pappus configuration as in Fig. 8 (interpretation of Farey addition in terms of Pappus theorem was even featured on the cover of a recent Intelligencer schwartz ).
Figure 8: Non-stretchable arrangement of $9$ projective pseudolines
Figure 9: Non-stretchable simple arrangement of $9$ affine pseudolines
Our main concern is, however, a slightly less complicated object, simple arrangements of $n$ affine pseudolines (no multiple intersection allowed). If $n\leq 7$, any one of them can be realized with lines. Yet, the configuration in Fig. 9 (found by L. Reeves and J. Wild again) with $8$ pseudolines cannot be straightened. A best current bound on number of simple arrangements of $n$ affine pseudolines was given by Justin Dallant dallant and is $2^{0.6496n^{2}}$. It is based on Zone Theorem (that will be discussed later). Compare this bound to a one given earlier for simple arrangements of $n$ lines.
The reason for introducing pseudolines is that their arrangement problems are easier and some aspects can be tackled solely combinatorially. we will witness this. It appears that many (but not all!) results about lines are more about their configuration that straightness property itself. There are other reasons to introduce pseudolines. For example, Suvorov suvorov gives two simple arrangements of $13$ lines in the affine plane. As pseudolines, one can be transformed into the other via a continuous deformation of the plane, so that the configuration always remains simple (flexible isotopy). This, however, cannot be done if curves are always required to remain straight (rigid isotopy).
Take the $\bigcup$ of $n$ $\triangle$ labelled $1$ through $n$. Go along the border of the $\bigcup$ and write down a number of $\triangle$ to which the next edge belongs. The result of this paper amounts to constructing a seq. of maximally possible length $\mathscr{U}(n)$ which serves as a model of such actual seq. The second step would be to show that this model is realisable in pseudoline geometry. Possibly, this is not too hard to show. The question whether a construction can further be straightened is of a very different sort. Stretchability problem is known to be NP-hard.
Lower Bound
It is straightforward to contrive $n\geq 1$ $\triangle$ whose $\bigcup$ is a simple polygon with 9n69n-6 edges. One only needs to choose $3$ points on $3$ different sides of the existing polygon, slightly protrude the points outwards, and construct a new $\triangle$ (Fig. 10A). The only two concerns are that no vertex of an initial polygon lie inside the newly constructed $\triangle$, and that the new polygon is simple. If one vertex of the initial polygon with reflex ($180^{\circ}$) angle lies inside a $\triangle$, we get an increase of edge count by $10$ (Fig. 10B). If two non-consecutive vertices with reflex angles are inside, the increase is by $11$ (Fig. 10C).
Figure 10: Adding one $\triangle$ changes an edge count.
With this idea in mind, let us take, as a start-up, three brown $\triangle$ as shown in Fig. 11. Their $\bigcup$ has $22$ edges. Consider a blue segment and two green circles as shown. The latter are inscribed into $\angle PXA$ and $\angle SYB$, each passing through points $A$ and $B$, respectively. Next, choose several points (labelled $1,2,\ldots,m=n-3$, left to right) strictly inside the blue segment. From every point $p$, draw “inner" tangents to both circles. Let the intersection with $PQ$ and $SR$ be, respectively, $p^{\prime}$ and $p^{\prime\prime}$. Fig. 11 shows example with $m=2$.
Figure 11: An inductive construction showing that $e(n)\geq 11n-11$ for $n\geq 4$.
We now miniaturelly protrude all points on $PQ$, SRSR and the blue segment to the exterior. Each of $m$ $\triangle$ adds $11$ edges. This gives the total count of 11n1111n-11.
In fact, nothing is specific about this start-up configuration. The construction works each time there exists an edge which is not adjacent to any vertex of an original $\triangle$ (so, $n\geq 3$). Take, for example, the green edge in Fig. 12 (the one with a control vertex numbered “47"). Its endpoints belong to a cyan and grey $\triangle$, respectively. Take a point $x$ inside this edge. We will construct a $\triangle$ whose vertex is $x$, only slightly protruded outwards. It will lie between black and brown vertices $a$ and $b$, respectively. Let left and right edges at these vertices intersect at point $p$ and $q$ (shown also on the same Fig. 12). Draw line $xp$ and $xq$ and find their intersections $y,z$ with the corresponding brown (shown in bold) and black segments on the boundary of the union. Slightly inflated $\triangle$ xyzxyz adds +11+11 in edge count. This mimics precisely the construction in Fig. 11. Also, vertices of triangles preserve a cyclic order.
Figure 12: $\bigcup$ of 88 $\triangle$ with $79$ edges, whose seq. is given by (1) (ignore $\triangle xyz$ for now). All vertices of 88 $\triangle$ are on the unit circle; few control vertices of a $79$-gon are labelled. $\triangle xyz$ is used in the inductive prove to show that $e(8+k)\geq 79+11k$.
Note that such lower bounds are situational. To see this, consider an Octastar in Fig. 5. Removing any 33 $\triangle$ give a collection of 55 $\triangle$ with at most $44$ edges in the $\bigcup$. In other words, we cannot expect to achieve maximum number of edges for $(n+1)$ $\triangle$ by finding the best addition to the existing best $n$ $\triangle$ configuration. Each $n$ should be treated individually, what we will do in the section ‘Maximal Sequences’.
Upper Bound
Take a regular $\bigcup$ of $n$ $\triangle$ with $M$ edges. For example, consider 88 $\triangle$ in Fig. 12. We start with a blue $\triangle$ (labelled $1$) and an edge just before the vertex with control label “10". All $\triangle$ are numbered $1$ through $8$. So, for example, black $\triangle$ gets number $4$. Now, go along the boundary in a positive direction and record what edges occur. This produces the cyclic seq. of length $79$:
$\underline{\underline{11}{\color[rgb]{1,0,0}3}\overline{2233{\color[rgb]{1,0,0}216}4455{\color[rgb]{1,0,0}4}66{\color[rgb]{1,0,0}418}7788{\color[rgb]{1,0,0}7}11}}\overline{{\color[rgb]{1,0,0}74}22}{\color[rgb]{1,0,0}4}3344{\color[rgb]{1,0,0}3276}$
$5566{\color[rgb]{1,0,0}5}77{\color[rgb]{1,0,0}521}88\underline{\underline{11{\color[rgb]{1,0,0}8}22{\color[rgb]{1,0,0}85}33{\color[rgb]{1,0,0}5}4455{\color[rgb]{1,0,0}4387}6677{\color[rgb]{1,0,0}6}88{\color[rgb]{1,0,0}63}}}.$ (1)
Take a fragment of the boundary contained between two vertices of the same blue $\triangle$. Say, the simply-underlined one in (1). Write down numbers of $\triangle$ whose RIGHT edges are encountered along the way. That is: ignore a number if its double has not yet occurred; otherwise, include it; from each doubled number, only one copy is included. This way the seq. $\mathcal{L}=123214546417871$ is obtained. What properties are intrinsic to it? First, no two adjacent numbers are equal. Second, though some of them may (and do) appear multiple times, such seq. do not contain a subseq. of the form abababab for any distinct pair $a,b\in\{1,2,\ldots,n\}$. This is a crucial property of regular $\triangle$ $\bigcup$. In the section ‘Lower envelopes’ we will see that as long as such seq. are allowed to have subseq. of the form abababab, their length can grow superlinearly.
How to prove the stated property? Since we go along the common boundary counter clock-wise, each leg has a direction. Drawing a full line through its endpoints, we can talk about right and left side of it. The same applies to any chosen edge. All the remaining edges split into lying to the right, to the left of it, or collinear. Also, for any two points on the same leg, we can say which one is behind and which one is ahead the other. Now, fix any number in the produced seq. of right edges. All the successive edges are either collinear (those with the same number), or lie to the right of it. Suppose, a subseq. abaaba belongs to such a seq. Let legs $a$ and $b$ intersect at a point $P$. In the process of recording right edges, we pass from leg $a$ to leg $b$, and then to leg $a$ again. Stop here. The point $P$ now lies behind. It is no longer possible to get beck to leg $b$, since all points of leg $b$ ahead of $P$ lie to the left of leg $a$.
Hence, $\mathcal{L}$ and all such seq. are exactly what is known to be a $(2,n)$-Davenport-Schinzel (or DS) seq. by definition sharir . The maximal length of $(2,n)$-DS is known to be 2n12n-1.
The proof of this fact is done by induction. Suppose, it is true for n1n-1. Consider a $(2,n)$-DS seq. $\mathcal{L}$. If not all symbols $1$ through $n$ appear, it is then a $(2,m)$-DS seq. for some $m<n$. Otherwise, reading left-to-right, assume that symbols appear in their natural order: $1,2,\ldots,n$ (one can always rename if needed). We claim that symbol $n$ appears in $\mathcal{L}$ only once. Assuming the opposite, let $\mathcal{L}$ contain a subseq. npnnpn for some $p\leq n-1$. Due to natural order convention, $\mathcal{L}$ must also contain a subseq. pnpnpnpn – a contradiction. Removing $n$ and concatenating two symbols which surround it in case they are equal, we get a (2,n1)(2,n-1)-DS seq. of a maximal length 2n32n-3. Thus, $\mathcal{L}$ has a length at most 2n3+22n-3+2. Note also that in case $(2,n)$-DS seq. is maximal, the last symbol to appear must always be surrounded by two identical symbols.
Let us return to our edge count. Take the same underlined fragment, but start from the other end. That is, write down numbers of $\triangle$ whose LEFT edge contribute to the boundary. This way, seq. $18781654613231$ is obtained. It is again a $(2,n)$-DS seq. Consequently, the total number of edges between two blue vertices is $\leq 4n-2$, with a stipulation that all blue edges were counted twice. We perform the same calculation for all three blue vertex pairs. Thus, if $\ell_{i}$ is the number of edges the $i$-th $\triangle$ contributes to the boundary of $M$-gon, then
$2\ell_{1}+\ell_{2}+\cdots+\ell_{n}\leq 12n-6.$ (2)
Summing over all $i$ yields $(n+1)M\leq n(12n-6)$. Thus, $M\leq 12n-18+\frac{18}{n+1}$. This looks great, with a caveat that in case $n=7$, the bound stands at $68$, while $67$ is the best we have yet managed to achieve $(45+11+11)$.
The inequality $e(n)<12n$ also implies the following. Take a regular $\bigcup$ of $n$ $\triangle$ with $e(n)$ edges. Choose $\triangle$ which contributes least to the common boundary. Its toll is thus $\leq 11$ edges. Remove this $\triangle$ from the $\bigcup$. Each edge not adjacent to its vertex reduces an edge count by $1$, while two edges at any vertex reduce this by $3$. Therefore, $e(n-1)\geq e(n)-14$.
The idea to separate left and right edges between two same-colored vertices was inspired by the proof of the aforementioned Zone Theorem given by Edelsbrunner at al. in edelsbrunner . In the latter paper though, ‘right’ and ‘left’ correspond to two directions along the same ray. Among several existing proofs of the Zone Theorem, it is the one which essentially employs $(2,n)$-DS sequences.
What this result claims? Consider $\mathcal{L}$, a simple arrangement of $n$ lines in the affine plane. The set $\mathbb{R}^{2}\setminus\mathcal{L}$ split into connected components called faces, which are polygons (in case face is bounded) or faces with two infinite edges. The size of a face is the number of its edges. Given a line $\ell$. If $\mathcal{L}\cup\{\ell\}$ is a simple arrangement, the zone of $\ell$ is a collection of all faces it crosses. The complexity of the zone of $\ell$ is the sum of sizes of all faces in the zone. This notion was first introduced by Chazelle at al. in zone , where an upper bound $6n$ for complexity was given. DS seq. approach yields 6n46n-4 (see edelsbrunner , p. 334, also sharir , p. 125). In horizon , Bern at al. strengthened this to $\lfloor 5.5n\rfloor-1$. Rom Pinchasi pinchasi later improved it to $\lfloor 5.5n\rfloor-3$ pinchasi . A construction in Fig. 13 (taken from horizon , just slightly enhanced) shows that the latter bound is almost optimal. Note the similarity with construction in Fig. 11.
Figure 13: An arrangement of $11$ lines; a complexity of the zone of blue line (colored in green) is $54=\lfloor 5.5\cdot 11\rfloor-6$; a size of each face in the zone is marked.
Maximal Sequences
The final insight comes from the observation that, depending on $n$, a certain amount of all $6n$ of the obtained seq. must be non-maximal. In terms of the upper bound, this only settles the case $n=7$, and reduces it by $1$ for $15\leq n\leq 18$. This seems a feeble contribution. However, such deeper analysis also gives an algorithm to construct a seq. of length $\mathscr{U}(n)$ which does not violate combinatorics of pseudolines.
For the starters, let us return to the example $\mathcal{L}_{1}=123214546417871$. Its length is $2\cdot 8-1$. So, it is maximal ($\mathbf{M}$ for short). Now, shift our attention to the cyan triangle. Or, better, the overlined fragment in (1). The produced seq. is also M. We argue that this extension to $\mathcal{L}_{1}$ can be reproduced from $\mathcal{L}_{1}$ itself. Indeed, starting from $\mathcal{L}_{1}$, remove all (but the last) $1$’s and place $2$ in the end. This results in $232454647871\ldots 2$. For this to be $\mathbf{M}$, dots should be replaced by two numbers: $\mathcal{L}_{2}=232454647871ab2$. We claim that $a,b$ can be uniquely reconstructed from a prefix $232454647871$ and the terminal number $2$. The argument behind this will now certainly look familiar. Let us methodically remove the last (reading left to right) symbol to appear, and merge its two neighbours. Such merge must always occur:
$2324546478\underline{71a}b2$ $\mathop{\longrightarrow}^{a=7}\,\,23245464\underline{787}b2\rightarrow$
$2324546\underline{47b}2$ $\mathop{\longrightarrow}^{b=4}\,\,232454642.$
This gives the only possible reconstruction: $\mathcal{L}_{2}=232454647871742$. The overlined fragment in (1) confirms it. In general, a completely identical argument proves the following statement.
Lemma.
Given a prefix $\mathcal{P}=1\ldots 2\ldots 3\ldots n$ and the terminal number $1$. If $\mathcal{P}1$ is a $(2,n)$-DS seq., there exists the unique $\mathcal{Q}$ such that $\mathcal{P}\mathcal{Q}1=\mathscr{S}(Q)$ is maximal.
This allows to iterate the procedure (reconstructed symbols are printed in red):
$\scriptstyle 123214546417871\rightarrow 232454647871{\color[rgb]{1,0,0}74}2\rightarrow 3454647871742{\color[rgb]{1,0,0}4}3\rightarrow$
$\scriptstyle 454647871742434\rightarrow 5678717234{\color[rgb]{1,0,0}3276}5\rightarrow 678717234327656\rightarrow$
$\scriptstyle 7871723432756{\color[rgb]{1,0,0}5}7\rightarrow 81234325657{\color[rgb]{1,0,0}521}8\rightarrow$
$\scriptstyle 123432565752181\rightarrow 2343256575281{\color[rgb]{1,0,0}8}2.$ (3)
This exactly (!) corresponds to actual seq. extracted from (1). So far, all of them are $\mathbf{M}$. We also notice that $\mathscr{S}^{9}(\mathcal{L})$ is obtained from $\mathcal{L}_{1}$ by increasing all terms by 1. As we will soon see, this is a general rule.
Let us now continue not with iterations of $\mathscr{S}$, but with actual sequences extracted from (1). For a while, $\mathbf{M}$ seq. continue. However, at a step $17$ (double underline), a nonmaximal sequence $12345436768{\color[rgb]{1,0,0}63}1$ (length $14$) occurs. Then it is followed with a $\mathbf{M}$ $2345436768631{\color[rgb]{1,0,0}3}2$. Next $2$ seq. are also $\mathbf{M}$. Then comes another shorther $56768612321645$ (length $14$), after that comes $\mathbf{M}$ $6768612321645{\color[rgb]{1,0,0}4}6$. Finally, after two more steps, everything closes up at the initial $1232145464178{\color[rgb]{1,0,0}7}1$. The conclusion: process almost always runs smoothly (and the next seq. is deterministic), only during $24$ steps two glitches occur. The total number of edges in the $\bigcup$ equals the number of red symbols in this sequence of sequences (each of length 15 or 14), plus $6\cdot 8$.
If the reader has already decided that this is too confusing to follow, this is true. Seeing such an arbitrary-looking collection of numbers in (3), it is impossible to fathom what is actually happening. In fact, with a help a visualisation of maximal $(2,n)$-sequences in terms of triangulations of regular $(n+1)$-gons (first discovered by D. Roselle roselle ; see also sharir , p.250-232), we will see that… almost nothing is happening!
To witness this, consider, as an example, a maximal $(2,9)$-DS sequence $12343536768632191$. First, divide it into $9$ decreasing portions as follows: | 1| 2| 3| 43| 53| 6| 76| 86321| 91||\,1|\,2\,|\,3\,|\,43\,|\,53\,|\,6\,|\,76\,|\,86321\,|\,91\,|. Given regular $(n+1)$-gon. For each portion starting from $i$, connect vertices with these particular labels to the vertex $i+1$. This way Fig. 14 is obtained.
Figure 14: A triangulation of a regular $10$-gon which corresponds to a maximal $(2,9)$-DS sequence $12343536768632191$.
It is straightforward to see that this not only gives the visualisation, but in fact a bijection between all triangulations of $(n+1)$-gon and essentially distinct maximal $(2,n)$-DS sequences (those which cannot be made equal after redefining or reordering initial set of symbols).
In our case, this triangulation convention is used for the natural order: $1,2,\ldots,n$. If the order is $i,i+1,\ldots i-1$, we adopt the following rule. First, rotate symbols cyclically, so that the order is $1,2,\ldots,n$. So, for $\mathscr{S}(\mathcal{L}_{1})=232454647871742$ one gets $121343536768631$. Now, use the standard convention for visualisation, only the distinguished yellow vertex should be moved to position $i$.
Let as apply this rule for visualising actual seq. extracted from (1). The result is presented as Fig. 15. It is surprising and relieving at the same time: nothing changes going from seq. 1 to seq. 2, except that the distinguished vertex moved one position. In light of such triangulation correspondence, Lemma become a tautology. Given a triangulation of a regular $n$-gon. Then it tells that if one removes all edges ending at one particular vertex, then there is only one way to complete the triangulation by adding edges starting from this particular vertex.
Such pattern continues up to seq. 8. Going to seq. 9 (it starts again at 1), the whole triangulation rotates one turn in the positive direction. This becomes clear if given some thought.
Now, what happens when a glitch (a shorter seq.) occurs? How it is succeeded by a $\mathbf{M}$ seq.? The same Fig. 15 (in the traditions of Intelligencer) might be considered as a proof without words of the following fact: at step $16\rightarrow 17$ (distinguished is vertex 8), one edge ending with $8$ is removed (at the same time, since the vertex changes from $8$ to $1$, a one-step rotation in positive direction occurs). At step $17\rightarrow 18$ (vertex 1) and edge with endpoint $1$ is added. After two more moves, another such incident takes place: one edge is removed, a new edge is reinstated during the next move.
We come to understand that the whole process runs as follows. Start from a distinguished vertex in a triangulation of a regular $(n+1)$-gon. Perform $3n$ moves. An $i$th move amounts to removing diagonals at this vertex; adding other diagonals (at the same vertex); counting their number $t_{i}$; switching to the next neighbour. After the whole operation, vertex n1n-1 is reached. The goal is to obtain the same triangulation, only rotated $3$ turns in the negative direction. The total combinatorial length of such $\triangle$ $\bigcup$ is then $\sum_{i=1}^{3n}t_{i}+6n$.
Figure 15: Triangulations corresponding to actual seq. in of RIGHT edges in Fig. 12; 1-9 also correspond exactly to seq. in (3)
If no deletions are necessary, it means that the triangulation coincides with itself if rotated by 3-3 moves. So, $\alpha(n)$ equals $6$ times the total number of edges in the triangulation, that is, 6(n+1+n2)=12n66(n+1+n-2)=12n-6, minus the sum of number of edges at three consecutive vertices. For $n+1=6$ and $9$ we see that this sum is equal to $9$ and $10$, respectively, independently of the start vertex. As expected, this gives $12\cdot 5-6-9=45$ and $12\cdot 8-6-10=80$ edges of the $\bigcup$.
Lower Envelopes
This section is very sketchy, mostly referring to the monograph of Sharir and Agarwal sharir . Take $7$ linear univariate functions defined on the real line. Label them $1$ through $7$. Let $f$ be defined as the minimum of these. If, while taking the minimal value, we jump from one line to the other, it is obviously no longer possible to go back to the abandoned one. Therefore, the support sequence (telling which function provides the minimum as one progresses along the $x$-axis) for $f$ consist of at most $7$ symbols.
However, if each linear function is defined on its own interval so that their $\bigcup$ is also an interval, the support sequence can be longer. This is clear. What is astonishing that it can be much, much, longer! As an example, consider $7$ segments as in Fig. 16.
Figure 16: The support for lower envelope of 77 segments with complexity $25$
The support seq. $1213145416164653737572767$ has length $25$. What properties does such seq. have? Similarly as with $\mathcal{L}$, no two adjacent symbols are equal. At the same time, we immediately notice its subseq. like $1313$, $1414$, $3535$, etc. This is a major distinction. We leave it as an exercise for th reader to prove that though subseq. abababab may be present, subseq. of the for ababaababa are not to be found. This is exactly the definition of a $(3,n)$-Davenport-Schintzel seq.
The foundational part of sharir (p. 12-42) is devoted to investigating $(3,n)$-DS seq. The story how their true complexity unravelled is inspiring. The final answer – the maximal length of such seq. is $2n\alpha(n)+O(n)$, a result by Seth Pettie pettie – tells little about the story. In the preface to the book is is recalled how Hart and Sharir went onto the discovery that superlinear grown is possible.
Here is another major achievement. Consider a $(3,n)$-DS, now of maximal possible length $27$ (see seq. A002004 in OEIS): $121314151616543272737475767$. It can be realized as a lower envelope of $7$ pseudosegments (connected parts of $7$ pseudolines). Can it be realized as a lower envelope of $7$ segments? In other words, can pseudosegment construction be stretched? I do not know. What is certain that there is little hope (and most likely, untrue; see sharir , p. 112) that a certain maximal $(3,n)$-DS seq. for large $n$ can serve as an envelope of $n$ segments. However, Wiernik and Sharir proved that there exist $(3,n)$-DS seq. whose length asymptotically differs by a factor of at most $2$ from the length of longest such seq. and which does serve as lower envelope of $n$ segments. Speaking philosophically, this shows that the Ackermann function, first discovered in logic and computability analysis, is intricately present in nature, if we assume that straight segments are. Along with Turing machines with record-halting time strauss , aperiodic hat monotile, tiles with high Heesch number, this is something what carries mathematics well outside the domain of natural sciences.
To finish the construction, consider Fig. 16 once again. If we imagine each segment completed to a triangle by two nearly vertical legs, we clearly can get a simple polygon. Thus, finding $i(n)$ rests on other fundamental problems about unions. The following problem might serve a good training field:
Problem.
Find $n$ for which $i(n)>e(n)$.
Outro
During the Journées Arithmétiques conference held in Vilnius, Lithuania in 2011, Artūras Dubickas gave a talk. Andrzej Schinzel, then in the audience, noted the question posed by the speaker. Later he wrote a note containing the solution (Fig. 17). That is how an autograph of a famous mathematician landed in a country-side of Northern Lithuania. Though it is a bit of a stretch, but numbers $n=1,2,5,8$ which give rise to perfect trangle unions are also in the Fibonacci sequence.
Figure 17: A handwritten note made by Andrzej Schinzel in Vilnius, 2011

### v4

@@
 Intro
-Latvian, Lithuanian and (now extinct) Old Prussian constitute a Baltic branch of Indoeuropean (IDE) languages.
-During a Baltic Culture Festival in Northern Lithuania, August 2024, the author fixed his eyes on a scene decoration, where Latvian national symbol Auseklis [^1] From Latvian aust, Lithuanian aušti, ‘to dawn’, which has common IDE roots with English East.
-was shown.
+Latvian, Lithuanian and (now extinct) Old Prussian constitute a Baltic branch of Indoeuropean languages.
+During a Baltic Culture Festival in Northern Lithuania, August 2024, the author fixed his eyes on a scene decoration depicting Latvian national symbol Auseklis [^1] From Latvian aust, Lithuanian aušti, ‘to dawn’, which is cognate with English East..
 It forms an octagonal star (Fig.
-1, left), which is just a standard octagram ($\{8,3\}$-gon in Schläfli notation).
-Though this cross is supposed to be drawn without lifting a stick from the ground (the reason behind this restriction is rooted, of course, in magic; same picture, left, where simple and double Lietuvēns crosses are shown), one clearly can distinguish $4$ triangles whose union forms a simple $16$-gon.
-With few more thoughts, it is immediate how one can increase edge count by slightly deforming all triangles.
+1, left), which is just a standard octagram ($\{8/3\}$-gon in Schläfli notation).
+Though this cross is supposed to be drawn without lifting a stick from the ground (the reason behind this restriction is rooted, undoubtedly, in magic; same picture, left, where simple and double ‘Lietuvēns’ crosses are shown [^2] For a background, see English Wikipedia page with this title.), one clearly can distinguish $4$ triangles whose union forms a simple $16$-gon.
+With few more thoughts, it is immediate how to increase an edge count by slightly deforming all triangles.
 It was therefore natural to wonder what is the best possible result.
 Such is a subjective genesis of the problem.
-And magic surely works!
+And the magic surely works!
 The author could not anticipate that hypostases of both crosses will eventually emerge in this purely mathematical endeavour.
@@
 Take $n$ $\triangle$.
-Its edges will henceforth be called legs, while edge will be reserved solely for a $\bigcup$ polygon.
+Its sides will henceforth be called legs, while edge will be solely reserved solely for a $\bigcup$.
 Any two of $\triangle$ intersect in at most $6$ points, and it is straightforward to contrive an example with $\sim 2n^{2}$ edges on the boundary of the $\bigcup$.
 However, in such construction many of the $\triangle$ appear to be very slim.
-If all $\triangle$ happen to have angles at least $\delta>0$ (so called $\delta$-fat"), the complement of the $\bigcup$ then has at most $O_{\delta}(n)$ connected components, and the boundary consists of $O_{\delta}(n\log\log n)$ edges fat .
+If all $\triangle$ happen to have angles at least $\delta>0$ (so called $\delta$-fat"), the complement of the $\bigcup$ has at most $O_{\delta}(n)$ connected components, and the boundary consists of $O_{\delta}(n\log\log n)$ edges fat .
 What happens, though, if the $\bigcup$ has no holes at all and its interior is connected?
 Let us formulate this precisely.
-Problem.
+Problems..
 Given an integer $n\geq 1$. Suppose, a simple polygon is a $\bigcup$ of $n$ $\triangle$. What is the maximal number of sides it can have? The following also applies.
-  • 1)
+  • $\mathrm{1)}$
     No further restrictions. This defines seq. $i(n)$.
-  • 2)
+  • $\mathrm{2)}$
     Vertices of $\triangle$ along the border of the $\bigcup$ are arranged in a cyclic way – seq. $e(n)$.
-  • 3)
+  • $\mathrm{3)}$
     Vertices of $\triangle$ are on the unit circle – seq. $c(n)$.
-  • 4)
-    Vertices are on the unit circle and are arranged in the cyclic way – seq. $r(n)$.
+  • $\mathrm{4)}$
+    Vertices are on the unit circle and are arranged in a cyclic way – seq. $r(n)$.
 It may seem surprising that the question “how many edges exactly?" has not been asked before.
@@
 Remove it.
-It is not difficult to see this reduces the edge count of the outer border by at most $\frac{i(n)}{n}+3$.
+It is not difficult to see that this reduces the edge count of the outer border by at most $\frac{i(n)}{n}+3$.
 Thus, as if
@@
 But a certain aspect of the phenomenon is captured.
-In the Section ‘Lower Envelopes’, employing the theorem of Micha Sharir Ady Wiernik (deep and unexpected result for any newcomer to the topic), we will see that one cannot expect to do much better than that: $i(n)=\Omega(n\alpha(n))$, where $\alpha(n)$ is the functional inverse of the so-called Ackermann function.
-It tends to $\infty$ with an argument, but extremely slowly.
+In the Section ‘Lower Envelopes’, employing the theorem of Micha Sharir & Ady Wiernik (deep and unexpected result for any newcomer to the topic), we will see that one cannot expect to do much better than that: $i(n)=\Omega(n\alpha(n))$, where $\alpha(n)$ is the functional inverse of the so-called (William) Ackermann function.
+There are several variants of the latter, and we do not give the precise definition.
+The one used here can be found in sharir , p.
+12-15.
+It is not essential which variant of the Ackermann function is meant, all asymptotic results hold.
+The function $\alpha$ tends to $\infty$ with an argument, but extremely slowly.
 Its value is $3$ in the interval $[5,16]$, and $4$ in the interval
 $\Bigg[17,\underbrace{2^{2^{\ldots 2^{2}}}}_{65536\text{ copies}}\,\Bigg].$
-It is impossible for a sane mind to fathom a place on the line of positive integers where this function surpasses $5$. There is a handful of other reasons why a plain answer is not to be expected.
+It is impossible for a sane mind to fathom a place on the positive half-line where this function surpasses $5$.
+There is a handful of other reasons why a plain answer about $i(n)$ is not to be expected.
+It is even worse with the seq.
+$c(n)$.
+Apart from $c(n)\leq i(n)$, we do not know where the things stand.
+No $\Omega$ result is currently available.
 Figure 2: I=mnmn+3I=mn-m-n+3 number of intersections of $n$-gon and $m$-gon for odd $m,n$ ($m=9$, $m=7$, $I=50$ here).
-However, the situation with cyclically arranged vertices is slightly different.
+However, the situation with cyclically arranged vertices is different.
 Such $\triangle$ $\bigcup$ will be called regular.
-The current paper, though mostly expository and derivative, introduces such objects for the first tiome and gives a new additional tool in dealing with $\bigcup$.
-Namely, triangulation shifts.
-Theoretically, an answer becomes within a reach.
-Sure, there exists problems in combinatorial geometry where an arm-stretch distance is still a considerable one.
-For example, it is known that two simple polygons with odd numbers ($m$ and $n$) of vertices can intersect in at least mnnm+3mn-n-m+3 points (Fig.
+By this it is also implied that the $\bigcup$ is proper – all vertices are among vertices of the $\bigcup$ polygon.
+In contrast from the situation with $i(n)$, regular $\bigcup$ is automatically a simple polygon.
+Henceforth, $e(n)$ will be our main focus.
+It now has a number A375986 in the OEIS oeis .
+The current paper, though mostly expository, introduces regular $\bigcup$ for the first time and gives a new tool to particularly deal with it.
+Theoretically at least, the precise answer for any $n$ becomes within a reach.
+Sure, there are problems in combinatorial geometry where an arm’s length distance is still a considerable one.
+For example, it is known that simple $m$-gon and $n$-gon with odd number of vertices can intersect in at least mnnm+3mn-n-m+3 points (Fig.
 2), and this is conjectured to be the peak value.
-Ackerman et al.
+Eyal Ackerman et al.
 ackerman proved that the maximum is indeed mnmn+Cmn-m-n+C.
-Accordingly, this is optimal, apart from the exact value of $C$ (still not rigorously known to be equal to $3$).
-Henceforth, we will concentrate on $e(n)$, referred here as the Politrigon seq..
-It now has a number A375986 in the OEIS oeis .
-What are its initial values?
+Accordingly, this is optimal, apart from the exact value of $C$, still not rigorously known to be equal to $3$.
+What are the initial values of $e(n)$?
 For $n=1$, there is nothing to prove.
 The $\bigcup$ of 22 $\triangle$, if a simple $M$-gon, can have any number of sides in the range $3\leq M\leq 10$, and also $M=12$.
-The latter is mostly familiar from a pleasingly symmetric hexagram known as the star of David.
+Only in the latter case the $\bigcup$ is regular, which is mostly familiar from a pleasingly symmetric hexagram known as the star of David.
 In order to prove that $e(3)=22$, case-by-case analysis can be performed.
-The example ($n=1,2,3$) is presented in Fig.
+The example is presented in Fig.
 10 (three brown $\triangle$).
-Figure 3: Regular $\bigcup$ of $4$ triangles with $33$ sides.
+Figure 3: Regular $\bigcup$ of 44 triangles with $33$ sides. This unique $33$-gon is calibrated to maximize the ratio of its area to a sum of squares of its edge-lengths.
 Figure 4: $10$ lines forming $25$ non-overlapping $\triangle$.
-Figure 5: All 44 perfect $\bigcup$ of triangles.
-These are exactly the configurations satisfying (n+1)e(n)=(12n6)n(n+1)e(n)=(12n-6)n.
-Allegorically speaking, the line of succession to the Eye of Providence continues with the Star of David, then with the Pentastar, and terminates with the Octastar.
-The last two have only rotational symmetry.
-Combinatorics of Pentastar cannot be reached from Octastar by removing 33 $\triangle$.
-This excludes any inductive construction of the best $\bigcup$.
-Moving onto $n=4$, similar computer-assisted overhaul show that $e(4)=33$ (Fig.
+Figure 5: All 33 perfect proper unions of triangles.
+These (along with a single $\triangle$) are exactly the unions satisfying (n+1)M=(12n6)n(n+1)M=(12n-6)n.
+The pentagram reincarnates as the Pentastar, and the Octastar is an unexpected hypostasis of Auseklis.
+Moving onto $n=4$, similar computer-assisted overhaul shows that $e(4)=33$ (Fig.
 3).
-If we knew all possible arrangements of $3n$ lines in the affine plane (seq.
-A090338), this catalogue would encode, among other things, the Politrigon seq.
-(and $i(n)$, too).
+If we knew all possible arrangements of $3n$ lines in general position in the affine plane (seq.
+A090338), this catalogue would encode, among other things, the seq.
+$e(n)$ (and $i(n)$, too).
 Such list (for $n$ lines) would also encode the solution to the Kobon $\triangle$ problem (seq.
@@
 In its full generality, Kobon’s problem is still unsolved.
-It is obvious that such full register of line configurations is of no use for large $n$ do to extreme growth (of order $2^{\Theta(n\log n)}$) of their total number.
+It is obvious that such a full register of line configurations is of no use for large $n$ do to extreme growth (of order $2^{\Theta(n\log n)}$) of their total number.
 In fact, the last exact value known is for $9$ lines, which amounts to $3.111.341$ simple arrangements.
 In our setting, this would settle, with some effort, the case $n=3$.
-Clearly, in all derivatives of line arrangement problem, other method should be looked for.
-And it does exist in regular $\triangle\bigcup$ case.
-With such tools at hand, it will be possible to demonstrate the unconditional inequality $(n+1)e(n)\leq(12n-6)n$.
-This leads to the discovery of two “perfect" $\triangle$ $\bigcup$ showing that $e(5)=45$ and $e(8)=80$ are possible.
-Together with a $\triangle$ and a hexagram, this is the complete list of maximally saturated $\triangle$ $\bigcup$.
-In other words, the only cases where the equality is achieved (Fig.
+Clearly, in any derivative of the line arrangement problem, other method should be looked for.
+And here is the main tool.
+Definition..
+Let $n\geq 2$.
+Given a (possibly, incomplete) triangulation $\mathcal{T}_{0}$ of a regular $(n+1)$-gon and a distinguished vertex $1$, other vertices labelled $2$ to $n+1$ in a positive direction.
+A $(n+1)$-triangulation shift is a collection of $3n+1$ triangulations $\mathcal{T}_{i}$ ($0\leq i\leq 3n$) on the same set of vertices.
+$\mathcal{T}_{i}$ is obtained from $\mathcal{T}_{i-1}$ by removal and a subsequent addition of some diagonals incident to a vertex $i$.
+The final triangulation $T_{3n}$ must coincide with $T_{0}$, but turned $3$ vertices clockwise.
+If $t_{i}$ is the number of diagonals of $\mathcal{T}_{i}$ incident to $i$, the weight of such a shift is defined to be $\omega(\mathcal{T}_{0})=\sum_{i=1}^{3n}t_{i}+6n$.
+So, we go round a regular $(n+1)$-gon exactly $3$ vertices short of $3$ full turns.
+It is allowed to slightly alter diagonals at any present vertex.
+The final result is the initial triangulation, only rotated by 3-3 turns.
+Intuitively, the biggest weight is achieved for those $\mathcal{T}_{0}$ which best coincide with itself after such a rotation.
+This object is of interest, but seems coming out of the blue.
+What’s its relation to $\triangle$ $\bigcup$?
+The answer will gradually unfold.
+Theorem..
+The seq.
+$e(n)$, $n\geq 1$, starts as $3$, $12$, $22$, $33$, $45$, $56$, $67$, $80$, $91$.
+Given any regular $\triangle$ $\bigcup$ with $M$ edges, there exists a $(n+1)$-triangulation shift with weight $M$.
+For $n\geq 5$,
+$e(n)\leq 11n-14+2\bigg\lfloor\frac{n+1}{3}\bigg\rfloor=\mathscr{U}(n),$
+and $\mathscr{U}(n)$ is the maximal weight among all $(n+1)$-triangulation shifts. Additionally, the inequality $11\leq e(n+1)-e(n)\leq 14$ holds.
+The last stated result gives a situational lower bound $e(n)\geq 11n-8$ ($n\geq 10$).
+This can always be improved, provided we manage to find a concrete $N\geq 10$ and a configuration of NN $\triangle$ with more than 11N811N-8 edges.
+In the course of the proof, we will frequently aberrate to cognate topics.
+Yet, the reader can be rest assured that there is no complicated scheme behind the proof.
+All steps will align into a transparent row.
+As an intermediate result, we will show that $(n+1)e(n)\leq(12n-6)n$.
+This a posteriori leads to the discovery of two “perfect" $\triangle$ $\bigcup$ and a demonstration that $e(5)=45$, $e(8)=80$.
+Together with a $\triangle$ and a hexagram, the four form the complete list of maximally saturated (that is, where the equality is achieved) $\triangle$ $\bigcup$ ( Fig.
 5).
-(exactly the very reincarnation of two magic crosses hinted to in the Introduction).
+This exactly corresponds to $(n+1)$-triangulation shifts where no diagonal is removed or added!
+In other words, to a triangulation which, if rotated by 3-3 vertices, coincides with itself.
 For $n\neq 1,2,5,8$, strict inequality must hold.
-Theorem..
-The seq. $e(n)$ starts as $3$, $12$, $22$, $33$, $45$, $56$, $67$, $80$, $91$. It also satisfies the following:
-$11\leq e(n+1)-e(n)\leq 14\quad(forn\geq 3),$
-$e(n)\leq 12n-18+I(10\leq n\leq 13)\quad(forn\geq 10).$
-The first inequality gives situational lower bound $e(n)\geq 11n-8$ ($n\geq 10$).
-This can always be improved, provided we manage to find a concrete $N\geq 10$ and a configuration of NN $\triangle$ with more than 11N811N-8 edges.
-The upper bound now stands in sharp contrast with a superlinear growth of the function $i(n)$.
-It not just bounds $e(n)$ from above, but in fact gives the exact combinatorial bound on first approximation.
-Without explaining this vague phrase now, here is a little trick to be played on the readers.
-Among all triangulations of a regular $n$-gon, choose the one whose maximal sum of number of diagonals adjacent at three consecutive vertices is minimal.
-Denote this minimum by $\delta(n)$ (see Fig.
-6).
-One of the many interpretations of Catalan numbers $C_{n}=\frac{1}{n+1}\binom{2n}{n}$ is that $C_{n-2}$ gives the number of triangulations of a convex $n$-gon.
-Since there are n3n-3 diagonals, $\delta(n)\geq\big\lceil 6-\frac{18}{n}\big\rceil$.
-The following statement is therefore a result of full check ($n\leq 17$) and explicit construction ($n\geq 18$):
-Proposition 1.
-The sequence $\delta(n)$, $n\geq 3$, starts as $0,2,3,3,4,5,4,5,5,5,5,5$, and then stabilizes at the value $6$ (see Fig. 6).
-The reader can verify that for $n\geq 2$, $\mathscr{U}(n)=12n-12-\delta(n+1)$ coincides with exact value of the seq.
-$e(n)$ and gives the exact upper combinatorial bound for $n\geq 10$.
-The relation of $\triangle$ $\bigcup$ to triangulations of $n$-gons is unexpected, and will be clarified later.
-But now, let us discuss what does it mean “combinatorial" vs.
+The asserted upper bound stands in sharp contrast with a superlinear growth of the function $i(n)$.
+The function $\mathscr{U}(n)$ can be thought as being the exact combinatorial bound on the first approximation.
+Let us now discuss what does it mean “combinatorial" vs.
 “geometric".
-[^3] Since we started from linguistics, it is apt to emphasize here that suffix dilemma ‘-ial’ vs.
-‘-ic’ essentially coincides with whether it is of Latin or of Greek origin.
+[^4] Quite a symbolic collision.
+On a linguistic level, the clash of suffixes ‘-ial’ vs.
+‘-ic’ is essentially equivalent to whether the root is of Latin or of Greek origin.
 In other words, what are pseudolines and what is stretchability.
-Figure 6: Triangulations of regular polygons which minimize the maximum of number of diagonals adjacent at three consecutive vertices. This minimum is denoted by $\alpha(n)$
 Pseudolines vs. Lines
 Apart from A006066 and A090338, there are other seq.
-in the OEIS made from kindred ingredients.
+in the OEIS made from kindred ingredients and emitting alike aroma.
 This means: seq., whose definition employs only well-familiar objects from Euclidean geometry, but whose calculation requisites rather delicate arrangements of these.
@@
 A250.001 (first proposed by Jonathan Wild), asks to find the number of ways to draw $n$ circles in the affine plane sloane ; oeis (definitions of admissible and equivalent configurations are intuitive and can be found in the references cited).
-In Figure 7, three (out of $173$) arrangements of $4$ circles are reproduced.
+In Fig.
+6, three (out of $173$) arrangements of $4$ circles are reproduced.
 Figure 6: Three ways to draw 4 circles. Configuration $\mathbf{c}$ has $12$ boundary arcs.
-Even having a ready example in front of the eyes, it takes some practice to correctly drag around centres and radii of four circles in order to achieve configurations $\mathbf{a}$ and $\mathbf{b}$.
+Even having a ready example in front of the eyes, it takes some practice to correctly drag with mouse centres and radii of four circles in order to achieve configurations $\mathbf{a}$ and $\mathbf{b}$.
 See a recent survey cutler for many more examples of such seq., associated with cutting of plane with all kinds of “exotic knives".
-On the other hand, pseudoline arrangements and stretchability problem are not only close in spirit, but play a significant role in determining $e(n)$.
+Going further, problems like pseudoline arrangements and stretchability are not only close in spirit and delicacy level, but play a significant role in determining $e(n)$.
 A pseudoline is a simple connected curve in the projective plane such that its removal does not disconnect $\mathbb{R}P^{2}$.
-Arrangement of pseudolines is a finite set of them so that each two meet exactly once, and there are at least two such intersection points (the affine case is modified accordingly).
+Arrangement of pseudolines is their finite set so that each two meet exactly once, and it is not a pencil (so, there must be at least two intersection points).
+The affine (Euclidean) case is modified accordingly.
 A survey by Felsner and Goodman felsner is another enriching reading.
 The most transparent arrangement of projective pseudolines that cannot be realized with lines (that is, “non-stretchable") is given by a twisted Pappus configuration as in Fig.
-8 (interpretation of Farey addition in terms of Pappus theorem was even featured on the cover of a recent Intelligencer schwartz ).
+7 (interpretation of Farey addition in terms of Pappus theorem was even featured on the cover of Intelligencer schwartz ).
 Figure 7: Non-stretchable arrangement of $9$ projective pseudolines
-Figure 9: Non-stretchable simple arrangement of $9$ affine pseudolines
-Our main concern is, however, a slightly less complicated object, simple arrangements of $n$ affine pseudolines (no multiple intersection allowed).
+Figure 8: One of 33 (out of $38612$) non-stretchable simple arrangement of $8$ affine pseudolines
+Our main concern is, however, a slightly less complicated object, simple arrangements of $n$ affine pseudolines (no multiple intersection allowed; seq.
+A090339).
+[^5] There are subtlenesses in what is being counted by any particular seq.
+If $n$ affine lines are in general position, there are no finite multiple intersection points, apart from that some subsets may be parallel.
+So, if treated as projective lines, such arrangement might not be simple.
+There are sequences which count arrangements with a cell distinguished and (un)oriented, etc.
 If $n\leq 7$, any one of them can be realized with lines.
@@
 Reeves and J.
-Wild again) with $8$ pseudolines cannot be straightened.
-A best current bound on number of simple arrangements of $n$ affine pseudolines was given by Justin Dallant dallant and is $2^{0.6496n^{2}}$.
-It is based on Zone Theorem (that will be discussed later).
-Compare this bound to a one given earlier for simple arrangements of $n$ lines.
+Wild) with $8$ pseudolines cannot be straightened.
+A best current bound on number of simple arrangements of $n$ affine pseudolines was given by Justin Dallant and is $2^{0.6496n^{2}}$ dallant .
+The proof is based on the Zone Theorem (that will be discussed later).
+Compare this bound to a one given earlier for simple arrangements of $n$ straight lines (as we will see, in the light of this discrepancy, the aforementioned result of Wiernik and Sharir becomes even more surprising).
 The reason for introducing pseudolines is that their arrangement problems are easier and some aspects can be tackled solely combinatorially.
-we will witness this.
-It appears that many (but not all!) results about lines are more about their configuration that straightness property itself.
+It appears that many (but not all!) results about lines are more about their configuration that straightness itself.
 There are other reasons to introduce pseudolines.
@@
 This, however, cannot be done if curves are always required to remain straight (rigid isotopy).
+See the end of Chapter 4 in felsner for few more spotlight distinctions.
 Take the $\bigcup$ of $n$ $\triangle$ labelled $1$ through $n$.
-Go along the border of the $\bigcup$ and write down a number of $\triangle$ to which the next edge belongs.
+Go along the border of the $\bigcup$ and write down a number of $\triangle$ to which the consequent edge belongs.
 The result of this paper amounts to constructing a seq.
-of maximally possible length $\mathscr{U}(n)$ which serves as a model of such actual seq.
+of maximal possible length $\mathscr{U}(n)$ which serves as a model of such actual seq.
 The second step would be to show that this model is realisable in pseudoline geometry.
-Possibly, this is not too hard to show.
+This might be not too hard to show.
 The question whether a construction can further be straightened is of a very different sort.
+The Section ‘Lower envelopes’ will contain an example of such a problem.
 Stretchability problem is known to be NP-hard.
@@
 It is straightforward to contrive $n\geq 1$ $\triangle$ whose $\bigcup$ is a simple polygon with 9n69n-6 edges.
-One only needs to choose $3$ points on $3$ different sides of the existing polygon, slightly protrude the points outwards, and construct a new $\triangle$ (Fig.
-10A).
+One only needs to choose $3$ points on $3$ different sides of the existing simple polygon, slightly protrude the points outwards, and construct a new $\triangle$ (Fig.
+9A).
+Figure 9: Adding one $\triangle$ changes an edge count.
 The only two concerns are that no vertex of an initial polygon lie inside the newly constructed $\triangle$, and that the new polygon is simple.
-If one vertex of the initial polygon with reflex ($180^{\circ}$) angle lies inside a $\triangle$, we get an increase of edge count by $10$ (Fig.
-10B).
-If two non-consecutive vertices with reflex angles are inside, the increase is by $11$ (Fig.
-10C).
-Figure 10: Adding one $\triangle$ changes an edge count.
+If, however, two incident edges are intersected by the same leg of a new $\triangle$, we get an increase of edge count by $10$ (Fig.
+9B).
+If two non-overlapping pairs of incident edges are intersected, the increase is by $11$ (Fig.
+9C).
 With this idea in mind, let us take, as a start-up, three brown $\triangle$ as shown in Fig.
-11.
+10.
 Their $\bigcup$ has $22$ edges.
@@
 Next, choose several points (labelled $1,2,\ldots,m=n-3$, left to right) strictly inside the blue segment.
-From every point $p$, draw “inner" tangents to both circles.
+From every such point $p$, draw inner tangents to both circles.
 Let the intersection with $PQ$ and $SR$ be, respectively, $p^{\prime}$ and $p^{\prime\prime}$.
@@
 Figure 10: An inductive construction showing that $e(n)\geq 11n-11$ for $n\geq 4$.
-We now miniaturelly protrude all points on $PQ$, SRSR and the blue segment to the exterior. Each of $m$ $\triangle$ adds $11$ edges. This gives the total count of 11n1111n-11.
+We now minutely protrude all points on $PQ$, SRSR and the blue segment to the exterior. Each of $m$ $\triangle$ adds $11$ edges. This gives the total count of 11n1111n-11.
 In fact, nothing is specific about this start-up configuration.
-The construction works each time there exists an edge which is not adjacent to any vertex of an original $\triangle$ (so, $n\geq 3$).
+The construction works each time there exists an edge of the $\bigcup$ which is not adjacent to any vertex of an original $\triangle$ (so, $n\geq 3$).
 Take, for example, the green edge in Fig.
@@
 Let left and right edges at these vertices intersect at point $p$ and $q$ (shown also on the same Fig.
-12).
-Draw line $xp$ and $xq$ and find their intersections $y,z$ with the corresponding brown (shown in bold) and black segments on the boundary of the union.
-Slightly inflated $\triangle$ xyzxyz adds +11+11 in edge count.
+11).
+Draw line $xp$ and $xq$ and find their intersections $y,z$ with the corresponding brown (shown in bold just a bit to the left of $p$) and black edges of the $\bigcup$.
+Slightly inflated $\triangle$ xyzxyz adds +11+11 in edge count, also vertices of $\triangle$ preserve a cyclic order.
 This mimics precisely the construction in Fig.
-11.
-Also, vertices of triangles preserve a cyclic order.
+10.
 Figure 11: $\bigcup$ of 88 $\triangle$ with $79$ edges, whose seq.
@@
 $\triangle xyz$ is used in the inductive prove to show that $e(8+k)\geq 79+11k$.
-Note that such lower bounds are situational.
-To see this, consider an Octastar in Fig.
-5.
-Removing any 33 $\triangle$ give a collection of 55 $\triangle$ with at most $44$ edges in the $\bigcup$.
-In other words, we cannot expect to achieve maximum number of edges for $(n+1)$ $\triangle$ by finding the best addition to the existing best $n$ $\triangle$ configuration.
-Each $n$ should be treated individually, what we will do in the section ‘Maximal Sequences’.
 Upper Bound
@@
 For example, consider 88 $\triangle$ in Fig.
-12.
-We start with a blue $\triangle$ (labelled $1$) and an edge just before the vertex with control label “10".
+11.
+We start with a blue $\triangle$ (labelled $1$) and an edge just before the vertex with a control label “10".
 All $\triangle$ are numbered $1$ through $8$.
-So, for example, black $\triangle$ gets number $4$.
-Now, go along the boundary in a positive direction and record what edges occur.
+For instance, black $\triangle$ gets number $4$.
+Now, go along the boundary in a positive direction and sequentially record what edge occurs.
 This produces the cyclic seq.
@@
 $\underline{\underline{11}{\color[rgb]{1,0,0}3}\overline{2233{\color[rgb]{1,0,0}216}4455{\color[rgb]{1,0,0}4}66{\color[rgb]{1,0,0}418}7788{\color[rgb]{1,0,0}7}11}}\overline{{\color[rgb]{1,0,0}74}22}{\color[rgb]{1,0,0}4}3344{\color[rgb]{1,0,0}3276}$
-$5566{\color[rgb]{1,0,0}5}77{\color[rgb]{1,0,0}521}88\underline{\underline{11{\color[rgb]{1,0,0}8}22{\color[rgb]{1,0,0}85}33{\color[rgb]{1,0,0}5}4455{\color[rgb]{1,0,0}4387}6677{\color[rgb]{1,0,0}6}88{\color[rgb]{1,0,0}63}}}.$ (1)
-Take a fragment of the boundary contained between two vertices of the same blue $\triangle$.
-Say, the simply-underlined one in (1).
+$5566{\color[rgb]{1,0,0}5}77{\color[rgb]{1,0,0}521}88\underline{\underline{11{\color[rgb]{1,0,0}8}22{\color[rgb]{1,0,0}85}33{\color[rgb]{1,0,0}5}4455{\color[rgb]{1,0,0}4387}6677{\color[rgb]{1,0,0}6}88{\color[rgb]{1,0,0}63}}}$ (1)
+Take separately a fragment of this boundary contained between two vertices of the same blue $\triangle$.
+Say, the singly-underlined one in (1).
 Write down numbers of $\triangle$ whose RIGHT edges are encountered along the way.
-That is: ignore a number if its double has not yet occurred; otherwise, include it; from each doubled number, only one copy is included.
+That is: ignore a number if its double has not yet occurred; otherwise, record it; from each doubled number, only one copy is included.
 This way the seq.
@@
 First, no two adjacent numbers are equal.
-Second, though some of them may (and do) appear multiple times, such seq.
+Second, though some numbers may (and do) appear multiple times, such seq.
 do not contain a subseq.
 of the form abababab for any distinct pair $a,b\in\{1,2,\ldots,n\}$.
-This is a crucial property of regular $\triangle$ $\bigcup$.
+This is a crucial attribute of a regular $\triangle$ $\bigcup$.
 In the section ‘Lower envelopes’ we will see that as long as such seq.
@@
 How to prove the stated property?
-Since we go along the common boundary counter clock-wise, each leg has a direction.
-Drawing a full line through its endpoints, we can talk about right and left side of it.
+Since we go along the boundary counter clockwise, each leg has a direction.
+Drawing a full line through its endpoints, we can talk about right and left sides of it.
 The same applies to any chosen edge.
-All the remaining edges split into lying to the right, to the left of it, or collinear.
+All the remaining edges split into lying to the right, to the left, or collinear with it.
 Also, for any two points on the same leg, we can say which one is behind and which one is ahead the other.
 Now, fix any number in the produced seq.
-of right edges.
+(like $\mathcal{L}$) of right edges.
 All the successive edges are either collinear (those with the same number), or lie to the right of it.
@@
 The point $P$ now lies behind.
-It is no longer possible to get beck to leg $b$, since all points of leg $b$ ahead of $P$ lie to the left of leg $a$.
-Hence, $\mathcal{L}$ and all such seq. are exactly what is known to be a $(2,n)$-Davenport-Schinzel (or DS) seq. by definition sharir . The maximal length of $(2,n)$-DS is known to be 2n12n-1.
-The proof of this fact is done by induction.
+It is no longer possible to get back unto $b$, since all points of leg $b$ ahead of $P$ lie to the left of leg $a$.
+Hence, $\mathcal{L}$ and all seq.
+of right edges between two same-colored vertices of a fixed $\triangle$ are what is known to be a $(2,n)$-Davenport-Schinzel (or DS) seq.
+by very definition sharir .
+The maximal length of $(2,n)$-DS is known to be 2n12n-1.
+The proof is done by induction.
 Suppose, it is true for n1n-1.
 Consider a $(2,n)$-DS seq.
-$\mathcal{L}$.
+$\mathcal{J}$.
 If not all symbols $1$ through $n$ appear, it is then a $(2,m)$-DS seq.
@@
 Otherwise, reading left-to-right, assume that symbols appear in their natural order: $1,2,\ldots,n$ (one can always rename if needed).
-We claim that symbol $n$ appears in $\mathcal{L}$ only once.
-Assuming the opposite, let $\mathcal{L}$ contain a subseq.
+We claim that symbol $n$ appears in $\mathcal{J}$ only once.
+Assuming the opposite, let $\mathcal{J}$ contain a subseq.
 npnnpn for some $p\leq n-1$.
-Due to natural order convention, $\mathcal{L}$ must also contain a subseq.
+Due to a natural order convention, $\mathcal{J}$ must also contain a subseq.
 pnpnpnpn – a contradiction.
@@
 of a maximal length 2n32n-3.
-Thus, $\mathcal{L}$ has a length at most 2n3+22n-3+2.
+Thus, $\mathcal{J}$ has a length at most 2n3+22n-3+2.
 Note also that in case $(2,n)$-DS seq.
@@
 It is again a $(2,n)$-DS seq.
-Consequently, the total number of edges between two blue vertices is $\leq 4n-2$, with a stipulation that all blue edges were counted twice.
+(non-maximal this time).
+Consequently, the total number of edges between two blue vertices is $\leq 4n-2$, with a stipulation that all blue ones were counted twice.
 We perform the same calculation for all three blue vertex pairs.
-Thus, if $\ell_{i}$ is the number of edges the $i$-th $\triangle$ contributes to the boundary of $M$-gon, then
+Thus, if $\ell_{i}$ is the number of edges the $i$-th $\triangle$ contributes to the boundary of a $M$-gon, then
 $2\ell_{1}+\ell_{2}+\cdots+\ell_{n}\leq 12n-6.$ (2)
@@
 Remove this $\triangle$ from the $\bigcup$.
-Each edge not adjacent to its vertex reduces an edge count by $1$, while two edges at any vertex reduce this by $3$.
+Each edge not adjacent to its vertex reduces an edge count by (at most) $1$, while two edges at any vertex reduce this number by (at most) $3$.
 Therefore, $e(n-1)\geq e(n)-14$.
@@
 In the latter paper though, ‘right’ and ‘left’ correspond to two directions along the same ray.
+Figure 12: An arrangement of $11$ lines; a complexity of the zone (colored in green) of a blue line is $54=\lfloor 5.5\cdot 11\rfloor-6$; a size of each face in the zone is marked.
 Among several existing proofs of the Zone Theorem, it is the one which essentially employs $(2,n)$-DS sequences.
-What this result claims?
-Consider $\mathcal{L}$, a simple arrangement of $n$ lines in the affine plane.
-The set $\mathbb{R}^{2}\setminus\mathcal{L}$ split into connected components called faces, which are polygons (in case face is bounded) or faces with two infinite edges.
+What it claims?
+Consider $\mathcal{K}$, a simple arrangement of $n$ lines in the affine plane.
+The set $\mathbb{R}^{2}\setminus\mathcal{K}$ split into connected components called faces, which are polygons (in case face is bounded) or faces with two infinite edges.
 The size of a face is the number of its edges.
 Given a line $\ell$.
-If $\mathcal{L}\cup\{\ell\}$ is a simple arrangement, the zone of $\ell$ is a collection of all faces it crosses.
+If $\mathcal{K}\cup\{\ell\}$ is a simple arrangement, the zone of $\ell$ is a collection of all faces it crosses.
 The complexity of the zone of $\ell$ is the sum of sizes of all faces in the zone.
 This notion was first introduced by Chazelle at al.
-in zone , where an upper bound $6n$ for complexity was given.
+in zone , where an upper bound $6n$ for the complexity was given.
 DS seq.
@@
 strengthened this to $\lfloor 5.5n\rfloor-1$.
-Rom Pinchasi pinchasi later improved it to $\lfloor 5.5n\rfloor-3$ pinchasi .
+Rom Pinchasi pinchasi later improved this to $\lfloor 5.5n\rfloor-3$ pinchasi .
 A construction in Fig.
 12 (taken from horizon , just slightly enhanced) shows that the latter bound is almost optimal.
-Note the similarity with construction in Fig.
-11.
-Figure 13: An arrangement of $11$ lines; a complexity of the zone of blue line (colored in green) is $54=\lfloor 5.5\cdot 11\rfloor-6$; a size of each face in the zone is marked.
+Note the similarity with the construction in Fig.
+10 (and also Fig.
+18 further in the text).
 Maximal Sequences
-The final insight comes from the observation that, depending on $n$, a certain amount of all $6n$ of the obtained seq.
+The final refinement in the upper bound comes from the observation that, depending on $n$, a certain amount of all $6n$ of the obtained seq.
 must be non-maximal.
-In terms of the upper bound, this only settles the case $n=7$, and reduces it by $1$ for $15\leq n\leq 18$.
-This seems a feeble contribution.
-However, such deeper analysis also gives an algorithm to construct a seq.
-of length $\mathscr{U}(n)$ which does not violate combinatorics of pseudolines.
-For the starters, let us return to the example $\mathcal{L}_{1}=123214546417871$.
-Its length is $2\cdot 8-1$.
-So, it is maximal ($\mathbf{M}$ for short).
+Such analysis also gives an algorithm to actually construct a seq.
+of length $\mathscr{U}(n)$ which does not violate combinatorics of a regular union of $n$ $\triangle$ whose sides are segments of pseudolines.
+For the starters, let us return to the example $\mathcal{L}=123214546417871$.
+It is maximal ($\mathbf{M}$ for short).
 Now, shift our attention to the cyan triangle.
@@
 is also M.
-We argue that this extension to $\mathcal{L}_{1}$ can be reproduced from $\mathcal{L}_{1}$ itself.
-Indeed, starting from $\mathcal{L}_{1}$, remove all (but the last) $1$’s and place $2$ in the end.
+We argue that this extension to $\mathcal{L}$ can be reproduced from $\mathcal{L}$ itself.
+Indeed, starting from $\mathcal{L}$, remove all (but the last) $1$’s, and place $2$ in the end.
 This results in $232454647871\ldots 2$.
 For this to be $\mathbf{M}$, dots should be replaced by two numbers: $\mathcal{L}_{2}=232454647871ab2$.
-We claim that $a,b$ can be uniquely reconstructed from a prefix $232454647871$ and the terminal number $2$.
-The argument behind this will now certainly look familiar.
-Let us methodically remove the last (reading left to right) symbol to appear, and merge its two neighbours.
+We assert that $a,b$ can be uniquely reconstructed from a prefix $232454647871$ and the terminal number $2$.
+The reasoning will now certainly look familiar.
+Let us methodically remove the last (reading left-to-right) symbol to appear, and merge its two neighbours.
 Such merge must always occur:
@@
 $2324546\underline{47b}2$ $\mathop{\longrightarrow}^{b=4}\,\,232454642.$
-This gives the only possible reconstruction: $\mathcal{L}_{2}=232454647871742$. The overlined fragment in (1) confirms it. In general, a completely identical argument proves the following statement.
-Lemma.
-Given a prefix $\mathcal{P}=1\ldots 2\ldots 3\ldots n$ and the terminal number $1$.
-If $\mathcal{P}1$ is a $(2,n)$-DS seq., there exists the unique $\mathcal{Q}$ such that $\mathcal{P}\mathcal{Q}1=\mathscr{S}(Q)$ is maximal.
-This allows to iterate the procedure (reconstructed symbols are printed in red):
+This gives the only possible reconstruction: $\mathscr{S}(\mathcal{L})=\mathcal{L}_{2}=232454647871742$.
+The overlined fragment in (1) confirms it.
+Let us iterate this procedure (reconstructed symbols are printed in red):
 $\scriptstyle 123214546417871\rightarrow 232454647871{\color[rgb]{1,0,0}74}2\rightarrow 3454647871742{\color[rgb]{1,0,0}4}3\rightarrow$
@@
 So far, all of them are $\mathbf{M}$.
-We also notice that $\mathscr{S}^{9}(\mathcal{L})$ is obtained from $\mathcal{L}_{1}$ by increasing all terms by 1.
-As we will soon see, this is a general rule.
-Let us now continue not with iterations of $\mathscr{S}$, but with actual sequences extracted from (1).
+We also notice that $\mathscr{S}^{9}(\mathcal{L})$ is obtained from $\mathcal{L}$ by increasing all terms by $1$.
+This turns out to be a general rule.
+Let us now continue not with iterations of $\mathscr{S}$, but with actual sequences drawn from (1).
 For a while, $\mathbf{M}$ seq.
 continue.
-However, at a step $17$ (double underline), a nonmaximal sequence $12345436768{\color[rgb]{1,0,0}63}1$ (length $14$) occurs.
+However, at a step $17$ (doubly underlined), a non-M sequence $12345436768{\color[rgb]{1,0,0}63}1$ (length $14$) occurs.
 Then it is followed with a $\mathbf{M}$ $2345436768631{\color[rgb]{1,0,0}3}2$.
@@
 Finally, after two more steps, everything closes up at the initial $1232145464178{\color[rgb]{1,0,0}7}1$.
-The conclusion: process almost always runs smoothly (and the next seq.
+The conclusion: process almost always runs smoothly (and so the next seq.
 is deterministic), only during $24$ steps two glitches occur.
-The total number of edges in the $\bigcup$ equals the number of red symbols in this sequence of sequences (each of length 15 or 14), plus $6\cdot 8$.
-If the reader has already decided that this is too confusing to follow, this is true.
+The total number of edges in the $\bigcup$ equals the number of red symbols added, plus $6\cdot 8$ (the amount of black symbols).
+If the reader has already decided that this is too confusing to follow, this is more than true.
 Seeing such an arbitrary-looking collection of numbers in (3), it is impossible to fathom what is actually happening.
-In fact, with a help a visualisation of maximal $(2,n)$-sequences in terms of triangulations of regular $(n+1)$-gons (first discovered by D.
+In fact, with a help a visualisation of maximal $(2,n)$-sequences in terms of triangulations of a regular $(n+1)$-gons (first discovered by D.
 Roselle roselle ; see also sharir , p.250-232), we will see that… almost nothing is happening!
-To witness this, consider, as an example, a maximal $(2,9)$-DS sequence $12343536768632191$.
+To witness, consider, as an example, a maximal $(2,9)$-DS sequence $12343536768632191$.
 First, divide it into $9$ decreasing portions as follows: | 1| 2| 3| 43| 53| 6| 76| 86321| 91||\,1|\,2\,|\,3\,|\,43\,|\,53\,|\,6\,|\,76\,|\,86321\,|\,91\,|.
 Given regular $(n+1)$-gon.
-For each portion starting from $i$, connect vertices with these particular labels to the vertex $i+1$.
+For each vertex $i+1$, connect it to all vertices in a portion which starts from $i$.
 This way Fig.
 13 is obtained.
+It is straightforward to see that this not only gives the visualisation, but in fact a bijection between all triangulations of $(n+1)$-gon with a marked vertex and essentially distinct maximal $(2,n)$-DS sequences (those which cannot be made equal after renaming or reordering initial set of symbols).
 Figure 13: A triangulation of a regular $10$-gon which corresponds to a maximal $(2,9)$-DS sequence $12343536768632191$.
-It is straightforward to see that this not only gives the visualisation, but in fact a bijection between all triangulations of $(n+1)$-gon and essentially distinct maximal $(2,n)$-DS sequences (those which cannot be made equal after redefining or reordering initial set of symbols).
-In our case, this triangulation convention is used for the natural order: $1,2,\ldots,n$.
-If the order is $i,i+1,\ldots i-1$, we adopt the following rule.
-First, rotate symbols cyclically, so that the order is $1,2,\ldots,n$.
-So, for $\mathscr{S}(\mathcal{L}_{1})=232454647871742$ one gets $121343536768631$.
-Now, use the standard convention for visualisation, only the distinguished yellow vertex should be moved to position $i$.
-Let as apply this rule for visualising actual seq.
-extracted from (1).
+Figure 14: Triangulations corresponding to actual seq. in of RIGHT edges in Fig. 11; 1-9 also correspond exactly to seq. in (3).
+Figure 15: Initial triangulations $\mathcal{T}_{0}$ of shifts with maximal weight $\omega(\mathcal{T}_{0})=12(n-1)-\gamma(n+1)$.
+In our case, this triangulation convention is used for the natural order $1,2,\ldots,n$.
+If the order is $i,i+1,\ldots,i-1$, we adopt the following rule.
+First, rotate symbols in the seq.
+cyclically, so that the natural order is restored.
+This way, for $\mathcal{L}_{2}=232454647871742$ one gets $121343536768631$.
+Now, use the standard convention for visualisation, only the distinguished yellow vertex should be moved to the position $i$.
+Let as apply this rule to grasp what actual seq.
+extracted from (1) embody.
 The result is presented as Fig.
-15.
-It is surprising and relieving at the same time: nothing changes going from seq.
-1 to seq.
-2, except that the distinguished vertex moved one position.
-In light of such triangulation correspondence, Lemma become a tautology.
-Given a triangulation of a regular $n$-gon.
-Then it tells that if one removes all edges ending at one particular vertex, then there is only one way to complete the triangulation by adding edges starting from this particular vertex.
-Such pattern continues up to seq. 8. Going to seq. 9 (it starts again at 1), the whole triangulation rotates one turn in the positive direction. This becomes clear if given some thought.
+14.
+It is simultaneously surprising and relieving: nothing changes going from seq.
+1 ($\mathcal{L}$) to seq.
+2 ($\mathcal{L}_{2}$), except that the distinguished vertex moved one position.
+In light of such triangulation correspondence, the fact that the map $\mathscr{S}$ is well-defined becomes a tautology.
+Given a full triangulation of a regular $(n+1)$-gon.
+Then, if all edges ending at one particular vertex are removed, there is only one way to complete the triangulation by adding edges starting from this very vertex.
+Sounds like a joke, though it is not.
+Moving on, such pattern (that is, the distinguished vertex jumps one ahead at a step) continues up to seq.
+8.
+Going to seq.
+9 (it starts again at 1), the whole triangulation rotates one turn in the positive direction, and the distinguished vertex jumps two vertices ahead.
+This becomes clear if given some thought.
 Now, what happens when a glitch (a shorter seq.) occurs?
@@
 At step $17\rightarrow 18$ (vertex 1) and edge with endpoint $1$ is added.
-After two more moves, another such incident takes place: one edge is removed, a new edge is reinstated during the next move.
+After two more steps, another such incident takes place: one edge is removed, a new edge is reinstalled in the next iteration.
+As a final acumen to understand the whole procedure, note the following.
+When a vertex $8$ is replaced by $1$, instead of a rotation and a jump by $2$ vertices ahead, we can imagine that no rotation occurs, and the jump is an ordinary one, just by $1$ vertex.
+This way, after $24$ steps we arrive at vertex $7$.
+From its perspective, we witness the same initial triangulation.
+Figure 16: The support for lower envelope of 77 segments with complexity $25$
+Figure 17: Geometrically perfect polygon $\bigcup$ of inequivalent combinatorics.
+Top: 55 hexagons (type 1).
+Bottom: 77 quadrileterals (type 2).
+The maximal ratio of edge-lengths is shown.
+This number is close to the optimal value.
 We come to understand that the whole process runs as follows.
-Start from a distinguished vertex in a triangulation of a regular $(n+1)$-gon.
+Start from a distinguished vertex $1$ in a triangulation of a regular $(n+1)$-gon.
 Perform $3n$ moves.
-An $i$th move amounts to removing diagonals at this vertex; adding other diagonals (at the same vertex); counting their number $t_{i}$; switching to the next neighbour.
-After the whole operation, vertex n1n-1 is reached.
-The goal is to obtain the same triangulation, only rotated $3$ turns in the negative direction.
+An $i$th move amounts to removing few diagonals at this vertex; adding some other diagonals at the same vertex; counting their total number $t_{i}$; passing to the neighbour ahead.
+The last vertex where the move is performed is n2n-2 (afterwards, the jump to n1n-1 is made, yet followed by a halt).
+The goal is to obtain the same triangulation, only rotated by 3-3 vertices.
 The total combinatorial length of such $\triangle$ $\bigcup$ is then $\sum_{i=1}^{3n}t_{i}+6n$.
-Figure 15: Triangulations corresponding to actual seq. in of RIGHT edges in Fig. 12; 1-9 also correspond exactly to seq. in (3)
-If no deletions are necessary, it means that the triangulation coincides with itself if rotated by 3-3 moves.
-So, $\alpha(n)$ equals $6$ times the total number of edges in the triangulation, that is, 6(n+1+n2)=12n66(n+1+n-2)=12n-6, minus the sum of number of edges at three consecutive vertices.
-For $n+1=6$ and $9$ we see that this sum is equal to $9$ and $10$, respectively, independently of the start vertex.
-As expected, this gives $12\cdot 5-6-9=45$ and $12\cdot 8-6-10=80$ edges of the $\bigcup$.
+But this is exactly the definition of a triangulation shift!
+If no deletions are necessary, the triangulation coincides with itself rotated by 3-3 moves.
+The only two non-trivial examples exist for $n+1=6$ and $n+1=9$ (see Fig.
+15).
+These exactly correspond to the Pentastar and the Octastar.
+For a generic $n+1$, if no deletions were necessary and we were to make the full $3$ turns, the weight would be 12n1212n-12.
+Define an auxiliary function $\gamma(\mathcal{T}_{0})=12n-12-\omega(\mathcal{T}_{0})$.
+It measures, so to say, the defect of a triangulation.
+It approximately equals the number of diagonals not shared by a 3-3 rotation of $\mathcal{T}_{0}$, plus a number of diagonals incident to some three consecutive vertices.
+That is, $\gamma(n)\sim n/3$.
+To precisely determine $\gamma$, we employ Python to do brute force calculations for $5\leq n\leq 16$.
+Fig.
+15 shows initial values and champions (starting from $n+1=7$, not unique) for the corresponding values of $\gamma$.
+After the pattern is found, it is a standard routine to demonstrate that for $n\geq 5$, $\gamma(n+1)=n+2-2\lfloor\frac{n+1}{3}\rfloor$.
 Lower Envelopes
-This section is very sketchy, mostly referring to the monograph of Sharir and Agarwal sharir .
 Take $7$ linear univariate functions defined on the real line.
@@
 This is clear.
-What is astonishing that it can be much, much, longer!
+What is astonishing that it can be much, much, longer.
 As an example, consider $7$ segments as in Fig.
 16.
-Figure 16: The support for lower envelope of 77 segments with complexity $25$
 The support seq.
@@
 At the same time, we immediately notice its subseq.
-like $1313$, $1414$, $3535$, etc.
+like $1313$, $1414$, $3535$, $4646$, etc.
 This is a major distinction.
-We leave it as an exercise for th reader to prove that though subseq.
+We leave it as an exercise for the reader to prove that though subseq.
 abababab may be present, subseq.
@@
 This is exactly the definition of a $(3,n)$-Davenport-Schintzel seq.
-The foundational part of sharir (p.
+The foundational part of the monograph by Sharir and Agarwal (sharir , p.
 12-42) is devoted to investigating $(3,n)$-DS seq.
@@
 is $2n\alpha(n)+O(n)$, a result by Seth Pettie pettie – tells little about the story.
-In the preface to the book is is recalled how Hart and Sharir went onto the discovery that superlinear grown is possible.
-Here is another major achievement.
-Consider a $(3,n)$-DS, now of maximal possible length $27$ (see seq.
+In the preface to the book it is recalled how Hart and Sharir went onto the discovery that superlinear grown is possible.
+Here is another major result.
+Consider a $(3,7)$-DS, now of maximal possible length $27$ (see seq.
 A002004 in OEIS): $121314151616543272737475767$.
@@
 112) that a certain maximal $(3,n)$-DS seq.
-for large $n$ can serve as an envelope of $n$ segments.
+for any sufficiently large $n$ can serve as a lower envelope of $n$ segments.
 However, Wiernik and Sharir proved that there exist $(3,n)$-DS seq.
-whose length asymptotically differs by a factor of at most $2$ from the length of longest such seq.
+whose length asymptotically differs by a factor of at most $2$ from the length of the longest such seq.
 and which does serve as lower envelope of $n$ segments.
 Speaking philosophically, this shows that the Ackermann function, first discovered in logic and computability analysis, is intricately present in nature, if we assume that straight segments are.
-Along with Turing machines with record-halting time strauss , aperiodic hat monotile, tiles with high Heesch number, this is something what carries mathematics well outside the domain of natural sciences.
+Along with Turing machines with record-halting time strauss , aperiodic hat monotile, tiles with high Heesch number, this is what carries mathematics well outside the domain of natural sciences.
 To finish the construction, consider Fig.
 16 once again.
-If we imagine each segment completed to a triangle by two nearly vertical legs, we clearly can get a simple polygon.
-Thus, finding $i(n)$ rests on other fundamental problems about unions.
-The following problem might serve a good training field:
-Problem.
+If we imagine each segment completed to a triangle by a left leg leaning slightly to the right and the right leg leaning slightly to the left, we clearly can get a simple polygon.
+Thus, finding $i(n)$ rests on other fundamental problems about $\bigcup$ and stretchability.
+Figure 18: A construction showing that $c(n)\geq 11n-15$ ($n=7$ in the example; improvement to 11n1311n-13 is immediate).
+Without the orange $\triangle$, the union becomes regular, and this construction yields $r(n)\geq 10n-10$ (also, 10n810n-8 can be achieved).
+Figure 19: A handwritten note made by Andrzej Schinzel in Vilnius, 2011.
+Open (but closable) problems
+Seq.
+$c(n)$ (vertices of all $\triangle$ are on the unit circle) and $r(n)$ (additionally, $\bigcup$ is regular) were not dealt with in this paper.
+Since $r(n)\leq e(n)$, this seq.
+is, obviously, bounded by a linear function.
+For $1\leq n\leq 8$ the equality holds.
+This might be the end of a friendship: numerical search produces cyclically arranged 99 $\triangle$ on the unit circle with $90$ edges, but no example with $91$ edges was found.
+On the other hand, Fig.
+18 (without an orange $\triangle$) gives a construction showing that $r(n)\geq 10n-10$.
+Two black curves are circle arcs.
+By pushing a red $\triangle$ towards a brown one and a violet one towards a blue, a reflex angle at vertex $v$ can be cut twice.
+This raises the lower bound to 10n810n-8.
+There is, however, some space for further improvement.
+Out of the four, the sequence $c(n)$ clearly stands out.
+Sure, $r(n)\leq c(n)\leq i(n)$.
+Fig.
+18 gives a construction showing that $c(n)\geq 11n-13$.
+Yet, at the moment it seems that nothing substantially more can be said.
+The lower envelope technique does not work here.
+Moreover, Agarwal and Sharir plainly pose a problem about the maximal complexity of the lower envelope of line segments whose both ends are anchored at a circle arc (sharir , p.
+112).
+Finally, triangulation shifts is a tool not specific to $\triangle$ $\bigcup$.
+It also helps to determine $e_{p}(n)$, a maximal number of edges a regular $\bigcup$ of $n$ $p$-gons can have.
+In the definition, just replace $3n$ moves with $pn$ moves, and a rotation by 3-3 with a rotation by p-p.
+Sure, such general problem feels now too messy and premature.
+Yet, here is its sub-problem of independent interest.
+A perfect proper $\bigcup$ of $n$ $p$-gons must satisfy $(n+1)e_{p}(n)=(4n-2)pn$.
+Each of them must correspond to a triangulation of a regular $(n+1)$-gon which, if rotated by p-p vertices, coincides with itself.
+This gives three basic types (in case $p=3$ represented by, respectively, the Star of David, the Pentastar, the Octastar).
+Type $1$: n=p1n=p-1, the number of edges M=2(p1)(2p3)M=2(p-1)(2p-3).
+Type $2$: n=2p1n=2p-1, M=(2p1)(4p3)M=(2p-1)(4p-3).
+Type $3$: n=3p1n=3p-1, M=2(2p1)(3p1)M=2(2p-1)(3p-1).
+If such perfect union can be realized with straight lines and, moreover, has a $n$-fold rotational symmetry, we call it geometrically perfect.
+For $p=6$, $n=5$ (type 1), and for $p=4$, $n=7$ (type 2) there exist two combinatorially different perfect unions.
+This corresponds to that fact that there exist two inequivalent (with respect to an action of a dihedral group $\mathbb{D}_{6}$) triangulations of a regular hexagon, and also two inequivalent (with respect to $\mathbb{D}_{8}$) triangulations of a regular octagon with a point-symmetry.
+All four can be realized geometrically (Fig.
+17).
+However, in the last example, the shortest edge should be really short.
+It is plausible that there exists only a finite number of geometrically perfect proper polygon $\bigcup$.
+If so, this might be a queer addition to other finite catalogues of geometric objects ($5$ Platonic solids, $11$ Archimedean tilings, $15$ types of pentagonal tilings, and so on).
+To round up, the author is happy to pass a torch to any interested reader. The first four of these questions might be immediately jumped upon.
+  • 1)
     Find $n$ for which $i(n)>e(n)$.
+  • 2)
+    Prove rigorously that $r(9)=90$.
+  • 3)
+    Improve the lower bound for $r(n)$.
+  • 4)
+    Find a perfect $\bigcup$ which is not realisable as geometrically perfect.
+  • 5)
+    What arguments are in favour and against the possible linear growth of $c(n)$?
+  • 6)
+    Can e(n+1)e(n)=14e(n+1)-e(n)=14 occur?
+Since $\mathscr{U}(n+1)-\mathscr{U}(n)$ from $n=5$ onwards periodically repeats the pattern $11,11,13$, the affirmative answer to the last question is not ruled out.
 Outro
-During the Journées Arithmétiques conference held in Vilnius, Lithuania in 2011, Artūras Dubickas gave a talk.
-Andrzej Schinzel, then in the audience, noted the question posed by the speaker.
-Later he wrote a note containing the solution (Fig.
-17).
-That is how an autograph of a famous mathematician landed in a country-side of Northern Lithuania.
-Though it is a bit of a stretch, but numbers $n=1,2,5,8$ which give rise to perfect trangle unions are also in the Fibonacci sequence.
-Figure 17: A handwritten note made by Andrzej Schinzel in Vilnius, 2011
+During the Journées Arithmétiques 2011 conference held Vilnius, Lithuania, Artūras Dubickas discussed a certain problem with Andrzej Schinzel.
+As was his custom, Prof.
+Schinzel later returned with a handwritten note containing the solution (see Fig.
+19 containing the statement; afterwards comes the solution).
+Prof.
+Dubickas kindly presented this note to me.
+Though it is a bit of a stretch, but numbers $n=1,2,5,8$, giving rise to perfect triangle unions, are also in the Fibonacci sequence.